A 591 g ball strikes a wall at 14.7 m/s and
rebounds at 11.1 m/s. The ball is in contact
with the wall for 0.045 s.
What is the magnitude of the average force
acting on the ball during the collision?
Answer in units of N.
Don't round answer.
1
Expert's answer
2019-12-04T08:38:53-0500

The impulse of a ball


"\Delta p=mv_f-mv_i"

"=0.591\:\rm kg\times( 11.1\: m/s-(-14.7\: m/s))=15.25\: N\cdot s"

The magnitude of the average force acting on the ball during the collision


"F=\frac{\Delta p}{\Delta t}=\frac{15.25\:\rm N\cdot s}{0.045\:\rm s}=339\:\rm N"